OpenStudy (anonymous):

MEDAL AND FAN Solve for x: 3/4 x + 5/8 = 4x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (***[isuru]***):

\[\frac{ 3 }{ 4 }x + \frac{ 5 }{ 8 } =4x\] multiply every term by 8 \[6x + 5 = 32x\] \[26x = 5\]\[x = \frac{ 5 }{26 }\] hope this will help ya!!

OpenStudy (anonymous):

So x = 5/26 ?